# Structure-Preserving Oscillation-Eliminating Discontinuous Galerkin Schemes for Ideal MHD Equations: Locally Divergence-Free and Positivity-Preserving

### v1

@@ 3 Positivity preservation for LDF OEDG method
$\mathcal{G}=\left\{\boldsymbol{U}=\left(\rho,\boldsymbol{m},\boldsymbol{B},E\right)^{\top}:\rho>0,~{}\mathcal{E}\left(\boldsymbol{U}\right):=E-\frac{1}{2}\left(\frac{{\boldsymbol{m}}^{2}}{\rho}+{\boldsymbol{B}}^{2}\right)>0\right\},$ (36)
where $\mathcal{E}\left(\boldsymbol{U}\right)$ denotes the internal energy. When the polynomial degree $k\geq 1$, method does not always ensure that $\boldsymbol{U}_{\sigma}^{n}\in\mathcal{G}$. Based on the Zhang–Shu framework [36, 37], to achieve a bound-preserving DG scheme for hyperbolic conservation laws, one must first identify a DG scheme that preserves the updated cell averages within the bounds. Subsequently, a local scaling bound-preserving limiter is used to ensure that the DG solution satisfies these bounds at the points of interest. According to the discoveries in [42, 44], for the conservative MHD system (1), the PP property of the updated cell averages of the DG solution is closely connected with a set of globally coupled discrete divergence-free conditions, which are, however, incompatible with the standard local scaling limiter. An effective way [43, 44] addressing this issue is to consider the following modified MHD system with an additional source term:
@@
where $\boldsymbol{S}\left(\boldsymbol{U}\right)=\left(0,\boldsymbol{B},\boldsymbol{u},\boldsymbol{u}\cdot\boldsymbol{B}\right)^{\top}$ and the right-hand side of (37) is termed the Godunov–Powell source term [48]. As demonstrated in [43, 44], this source term can help to eliminate the effect of divergence error on the PP property.
In this section, we will utilize the optimal convex decomposition approach proposed in [53, 54] to derive an efficient, provably PP LDF OEDG method. This method will be built on a PP HLL flux, a properly upwind discretization of the Godunov–Powell source term, and a simple scaling PP limiter. Our method automatically ensures that the updated cell averages of the OEDG solution remain within $\mathcal{G}$, and the PP limiter enforces the values of the OEDG solution in $\mathcal{G}$ at any points of interest.
@@ 3.1 Positivity-preserving HLL flux
$\alpha_{\ell,l}\left(\boldsymbol{U},\tilde{\boldsymbol{U}}\right):=\min\left\{u_{\ell},\frac{\sqrt{\rho}u_{\ell}+\sqrt{\tilde{\rho}}\tilde{u}_{\ell}}{\sqrt{\rho}+\sqrt{\tilde{\rho}}}\right\}-\mathscr{C}_{\ell}\left(\boldsymbol{U}\right)-\frac{|\boldsymbol{B}-\tilde{\boldsymbol{B}}|}{\sqrt{\rho}+\sqrt{\tilde{\rho}}},\quad\ell=1,2,$ (39)
$\alpha_{\ell,r}\left(\boldsymbol{U},\tilde{\boldsymbol{U}}\right):=\max\left\{u_{\ell},\frac{\sqrt{\rho}u_{\ell}+\sqrt{\tilde{\rho}}\tilde{u}_{\ell}}{\sqrt{\rho}+\sqrt{\tilde{\rho}}}\right\}+\mathscr{C}_{\ell}\left(\boldsymbol{U}\right)+\frac{|\boldsymbol{B}-\tilde{\boldsymbol{B}}|}{\sqrt{\rho}+\sqrt{\tilde{\rho}}},\quad\ell=1,2,$ (40)
and
$\alpha_{\ell,\star}\left(\boldsymbol{U},\tilde{\boldsymbol{U}}\right):=\max\left\{|u_{\ell}|,\left|\frac{\sqrt{\rho}u_{\ell}+\sqrt{\tilde{\rho}}\tilde{u}_{\ell}}{\sqrt{\rho}+\sqrt{\tilde{\rho}}}\right|\right\}+\mathscr{C}_{\ell}\left(\boldsymbol{U}\right)+\frac{|\boldsymbol{B}-\tilde{\boldsymbol{B}}|}{\sqrt{\rho}+\sqrt{\tilde{\rho}}},\quad\ell=1,2,$ (41)
@@
$\mathscr{C}_{\ell}\left(\boldsymbol{U}\right):=\frac{1}{\sqrt{2}}\left[\mathscr{C}_{s}^{2}+\frac{\boldsymbol{B}^{2}}{\rho}+\sqrt{\left({\mathscr{C}_{s}}^{2}+\frac{\boldsymbol{B}^{2}}{\rho}\right)^{2}-4\frac{\mathscr{C}_{s}^{2}{B_{\ell}}^{2}}{\rho}}\right]^{\frac{1}{2}}\quad\text{with}\quad\mathscr{C}_{s}=\sqrt{\frac{\left(\gamma-1\right)p}{2\rho}},\quad\ell=1,2.$ (42)
@@ 3.2 Upwind discretization of Godunov–Powell source term
the DDF condition is not necessary for PP property due to the application of properly discretized Godunov-Powell source term. In fact, properly discretized Godunov-Powell source term helps eliminate the effect of the jump of the normal magnetic field on the cell interface.
Since the PP property of the scheme is only related to the discrete equations satisfied by the cell averages, we only consider adding Godunov-Powell source term to the evolution equations of the cell averages. That is, only the evolution of the cell averages adopts discrete scheme of the symmetrizable MHD equations (37), while the evolutions of the higher-order “moments” still use discrete scheme of the conservative MHD equations, as shown in Section 2.
Let $\boldsymbol{U}_{\sigma}^{n}\left(x,y\right)$ denote the approximate polynomial solution of degree $k$ at $n$-th time level after LDF OE procedure, and $\bar{\boldsymbol{U}}_{ij}^{n}$ be the cell average over cell $I_{ij}$ at $n$-th time level. For one-order forward Euler method, the evolution equations satisfied by the cell averages in the symmetrizable MHD equations with the Godunov-Powell source term are given by
@@
$\boldsymbol{L}_{ij}\left(\boldsymbol{U}_{\sigma}^{n}\right)$ $:=-\frac{1}{\Delta x}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left[\left(\widehat{\boldsymbol{F}_{1}}_{,i+\frac{1}{2}}\left(y_{j}^{(\mu)}\right)-\widehat{\boldsymbol{F}_{1}}_{,i-\frac{1}{2}}\left(y_{j}^{(\mu)}\right)\right)\right.$ (46)
$+\left.\mathscr{B}_{1,i+\frac{1}{2}}\left(y_{j}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{\sigma}^{n}\left(x_{i+\frac{1}{2}}^{-},y_{j}^{(\mu)}\right)\right)-\mathscr{B}_{1,i-\frac{1}{2}}\left(y_{j}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{\sigma}^{n}\left(x_{i-\frac{1}{2}}^{+},y_{j}^{(\mu)}\right)\right)\right].$
$-\frac{1}{\Delta y}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left[\left(\widehat{\boldsymbol{F}_{2}}_{,j+\frac{1}{2}}\left(x_{i}^{(\mu)}\right)-\widehat{\boldsymbol{F}_{2}}_{,j-\frac{1}{2}}\left(x_{i}^{(\mu)}\right)\right)\right.$
$+\left.\mathscr{B}_{2,j+\frac{1}{2}}\left(x_{i}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{\sigma}^{n}\left(x_{i}^{(\mu)},y_{j+\frac{1}{2}}^{-}\right)\right)-\mathscr{B}_{2,j-\frac{1}{2}}\left(x_{i}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{\sigma}^{n}\left(x_{i}^{(\mu)},y_{j-\frac{1}{2}}^{+}\right)\right)\right].$
Here $\left\{x_{i}^{(\mu)}\right\}_{\mu=1}^{q}$ and $\left\{y_{j}^{(\mu)}\right\}_{\mu=1}^{q}$ are the $q$-point Gauss-Legendre quadrature nodes with $2q-1\geq 2k$ in $\left[x_{i-\frac{1}{2}},x_{i+\frac{1}{2}}\right]$ and $\left[y_{j-\frac{1}{2}},y_{j+\frac{1}{2}}\right]$, respectively, and $\left\{\omega_{\mu}^{G}\right\}_{\mu=1}^{q}$ are the associated weights. In (46),
$\widehat{\boldsymbol{F}_{1}}_{,i+\frac{1}{2}}\left(y\right):=\widehat{\boldsymbol{F}_{1}}\left(\boldsymbol{U}_{\sigma}^{n}\left(x_{i+\frac{1}{2}}^{-},y\right),\boldsymbol{U}_{\sigma}^{n}\left(x_{i+\frac{1}{2}}^{+},y\right)\right),\$
and
$\widehat{\boldsymbol{F}_{2}}_{,j+\frac{1}{2}}\left(x\right):=\widehat{\boldsymbol{F}_{2}}\left(\boldsymbol{U}_{\sigma}^{n}\left(x,y_{j+\frac{1}{2}}^{-}\right),\boldsymbol{U}_{\sigma}^{n}\left(x,y_{j+\frac{1}{2}}^{+}\right)\right),\$
where $\widehat{\boldsymbol{F}_{1}}$ and $\widehat{\boldsymbol{F}_{2}}$ denote HLL numerical fluxes. In (46), we take
$\mathscr{B}_{1,i\pm\frac{1}{2}}\left(y\right)=\frac{-\mathscr{V}_{1,i\pm\frac{1}{2}}^{\mp}\left(y\right)}{\mathscr{V}_{1,i\pm\frac{1}{2}}^{+}\left(y\right)-\mathscr{V}_{1,i\pm\frac{1}{2}}^{-}\left(y\right)}\left[{\left(B_{1}\right)}_{\sigma}^{n}\left(x_{i\pm\frac{1}{2}}^{+},y\right)-{\left(B_{1}\right)}_{\sigma}^{n}\left(x_{i\pm\frac{1}{2}}^{-},y\right)\right],$
and
$\mathscr{B}_{2,j\pm\frac{1}{2}}\left(x\right)=\frac{-\mathscr{V}_{2,j\pm\frac{1}{2}}^{\mp}\left(x\right)}{\mathscr{V}_{2,j\pm\frac{1}{2}}^{+}\left(x\right)-\mathscr{V}_{2,j\pm\frac{1}{2}}^{-}\left(x\right)}\left[{\left(B_{2}\right)}_{\sigma}^{n}\left(x,y_{j\pm\frac{1}{2}}^{+}\right)-{\left(B_{2}\right)}_{\sigma}^{n}\left(x,y_{j\pm\frac{1}{2}}^{-}\right)\right],$
and with the definitions of
$\mathscr{V}_{1,i+\frac{1}{2}}^{\pm}\left(y\right):=\mathscr{V}_{1}^{\pm}\left(\boldsymbol{U}_{\sigma}^{n}\left(x_{i+\frac{1}{2}}^{-},y\right),\boldsymbol{U}_{\sigma}^{n}\left(x_{i+\frac{1}{2}}^{+},y\right)\right)$
and
@@
where $\mathscr{V}_{1}^{\pm}$ and $\mathscr{V}_{2}^{\pm}$ are the wave speeds in the HLL flux, which are same as that in Section 2. It is worth mentioning that the discretized Godunov-Powell source term in (46) is motivated from the analysis for provably PP property, as shown in the following Theorem 2.
In fact, we can split the discrete divergence into two parts:
@@
$\text{div}_{I_{ij}}\boldsymbol{B}^{n}=\frac{1}{\Delta x\Delta y}\int_{I_{ij}}\nabla\cdot\boldsymbol{B}^{n}\left(\boldsymbol{x}\right)\mathrm{~{}d}\boldsymbol{x}=\frac{1}{\Delta x\Delta y}\oint_{I_{ij}}\boldsymbol{B}^{n}\left(\boldsymbol{x}\right)\cdot\boldsymbol{n}\mathrm{~{}d}l$
$=\frac{1}{\Delta x}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left[\frac{\mathscr{V}_{1,i+\frac{1}{2}}^{+}\left(y_{j}^{(\mu)}\right){\left(B_{1}\right)}_{\sigma}^{n}\left(x_{i+\frac{1}{2}}^{-},y_{j}^{(\mu)}\right)-\mathscr{V}_{1,i+\frac{1}{2}}^{-}\left(y_{j}^{(\mu)}\right){\left(B_{1}\right)}_{\sigma}^{n}\left(x_{i+\frac{1}{2}}^{+},y_{j}^{(\mu)}\right)}{\mathscr{V}_{1,i+\frac{1}{2}}^{+}\left(y_{j}^{(\mu)}\right)-\mathscr{V}_{1,i+\frac{1}{2}}^{-}\left(y_{j}^{(\mu)}\right)}\right.$
$+\left.\frac{\mathscr{V}_{1,i-\frac{1}{2}}^{+}\left(y_{j}^{(\mu)}\right){\left(B_{1}\right)}_{\sigma}^{n}\left(x_{i-\frac{1}{2}}^{+},y_{j}^{(\mu)}\right)-\mathscr{V}_{1,i-\frac{1}{2}}^{-}\left(y_{j}^{(\mu)}\right){\left(B_{1}\right)}_{\sigma}^{n}\left(x_{i-\frac{1}{2}}^{-},y_{j}^{(\mu)}\right)}{\mathscr{V}_{1,i-\frac{1}{2}}^{+}\left(y_{j}^{(\mu)}\right)-\mathscr{V}_{1,i-\frac{1}{2}}^{-}\left(y_{j}^{(\mu)}\right)}\right]$
$+\frac{1}{\Delta y}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left[\frac{\mathscr{V}_{2,j+\frac{1}{2}}^{+}\left(x_{i}^{(\mu)}\right){\left(B_{2}\right)}_{\sigma}^{n}\left(x_{i}^{(\mu)},y_{j+\frac{1}{2}}^{-}\right)-\mathscr{V}_{2,j+\frac{1}{2}}^{-}\left(x_{i}^{(\mu)}\right){\left(B_{2}\right)}_{\sigma}^{n}\left(x_{i}^{(\mu)},y_{j+\frac{1}{2}}^{+}\right)}{\mathscr{V}_{2,j+\frac{1}{2}}^{+}\left(x_{i}^{(\mu)}\right)-\mathscr{V}_{2,j+\frac{1}{2}}^{-}\left(x_{i}^{(\mu)}\right)}\right.$
$+\left.\frac{\mathscr{V}_{2,j-\frac{1}{2}}^{+}\left(x_{i}^{(\mu)}\right){\left(B_{2}\right)}_{\sigma}^{n}\left(x_{i}^{(\mu)},y_{j-\frac{1}{2}}^{+}\right)-\mathscr{V}_{2,j-\frac{1}{2}}^{-}\left(x_{i}^{(\mu)}\right){\left(B_{2}\right)}_{\sigma}^{n}\left(x_{i}^{(\mu)},y_{j-\frac{1}{2}}^{-}\right)}{\mathscr{V}_{2,j-\frac{1}{2}}^{+}\left(x_{i}^{(\mu)}\right)-\mathscr{V}_{2,j-\frac{1}{2}}^{-}\left(x_{i}^{(\mu)}\right)}\right]$
$=\text{div}_{\text{within}}\boldsymbol{B}^{n}+\text{div}_{\text{faces}}\boldsymbol{B}^{n},$
where
@@
$\text{div}_{\text{within}}\boldsymbol{B}^{n}=\frac{1}{\Delta x}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left[{\left(B_{1}\right)}_{\sigma}^{n}\left(x_{i+\frac{1}{2}}^{-},y_{j}^{(\mu)}\right)-{\left(B_{1}\right)}_{\sigma}^{n}\left(x_{i-\frac{1}{2}}^{+},y_{j}^{(\mu)}\right)\right]$
@@
$+\frac{1}{\Delta y}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left[{\left(B_{2}\right)}_{\sigma}^{n}\left(x_{i}^{(\mu)},y_{j+\frac{1}{2}}^{-}\right)-{\left(B_{2}\right)}_{\sigma}^{n}\left(x_{i}^{(\mu)},y_{j-\frac{1}{2}}^{+}\right)\right],$
and
@@
$\text{div}_{\text{faces}}\boldsymbol{B}^{n}=\frac{1}{\Delta x}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left[\frac{-\mathscr{V}_{1,i+\frac{1}{2}}^{-}\left(y_{j}^{(\mu)}\right)}{\mathscr{V}_{1,i+\frac{1}{2}}^{+}\left(y_{j}^{(\mu)}\right)-\mathscr{V}_{1,i+\frac{1}{2}}^{-}\left(y_{j}^{(\mu)}\right)}\left({\left(B_{1}\right)}_{\sigma}^{n}\left(x_{i+\frac{1}{2}}^{+},y_{j}^{(\mu)}\right)-{\left(B_{1}\right)}_{\sigma}^{n}\left(x_{i+\frac{1}{2}}^{-},y_{j}^{(\mu)}\right)\right)\right.$
$+\left.\frac{\mathscr{V}_{1,i-\frac{1}{2}}^{+}\left(y_{j}^{(\mu)}\right)}{\mathscr{V}_{1,i-\frac{1}{2}}^{+}\left(y_{j}^{(\mu)}\right)-\mathscr{V}_{1,i-\frac{1}{2}}^{-}\left(y_{j}^{(\mu)}\right)}\left({\left(B_{1}\right)}_{\sigma}^{n}\left(x_{i-\frac{1}{2}}^{+},y_{j}^{(\mu)}\right)-{\left(B_{1}\right)}_{\sigma}^{n}\left(x_{i-\frac{1}{2}}^{-},y_{j}^{(\mu)}\right)\right)\right]$

### v2

@@ 3 Positivity preservation for LDF OEDG method
$\mathcal{G}=\left\{\boldsymbol{U}=\left(\rho,\boldsymbol{m},\boldsymbol{B},E\right)^{\top}:\rho>0,~{}\mathcal{E}\left(\boldsymbol{U}\right):=E-\frac{1}{2}\left(\frac{{\|\boldsymbol{m}\|}^{2}}{\rho}+{\|\boldsymbol{B}\|}^{2}\right)>0\right\},$ (36)
where $\mathcal{E}\left(\boldsymbol{U}\right)$ denotes the internal energy. When the polynomial degree $k\geq 1$, the OEDG method does not always ensure that $\boldsymbol{U}_{\sigma}^{n}\in\mathcal{G}$. Based on the Zhang–Shu framework [39, 40], to achieve a bound-preserving DG scheme for hyperbolic conservation laws, one must first identify a DG scheme that preserves the updated cell averages within the bounds. Subsequently, a local scaling bound-preserving limiter is used to ensure that the DG solution satisfies these bounds at all the points of interest. According to discoveries in [46, 48], for the conservative MHD system (1), the PP property of the updated cell averages of the DG solution is closely connected with a set of globally coupled discrete divergence-free conditions, which, however, are incompatible with the standard local scaling limiter. An effective way [47, 48] to address this issue is to consider the following symmetrizable MHD system with an additional source term:
@@
where $\boldsymbol{S}\left(\boldsymbol{U}\right)=\left(0,\boldsymbol{B},\boldsymbol{u},\boldsymbol{u}\cdot\boldsymbol{B}\right)^{\top}$, and the right-hand side of (37) is termed the Godunov–Powell source term [51, 52]. As demonstrated in [47, 48], this source term can help to eliminate the effect of divergence error on the PP property.
In this section, we will utilize the optimal convex decomposition approach proposed in [58, 57] to derive an efficient, provably PP LDF OEDG method. This method will be built on a PP HLL flux, a properly upwind discretization of the Godunov–Powell source term, and a simple scaling PP limiter. Our method automatically ensures that the updated cell averages of the OEDG solution remain within $\mathcal{G}$, while the PP limiter enforces the values of the OEDG solution in $\mathcal{G}$ at any points of interest.
@@ 3.1 Positivity-preserving HLL flux
$\alpha_{\ell,l}\left(\boldsymbol{U},\tilde{\boldsymbol{U}}\right):=\min\left\{u_{\ell},\frac{\sqrt{\rho}u_{\ell}+\sqrt{\tilde{\rho}}\tilde{u}_{\ell}}{\sqrt{\rho}+\sqrt{\tilde{\rho}}}\right\}-\mathscr{C}_{\ell}\left(\boldsymbol{U}\right)-\frac{\|\boldsymbol{B}-\tilde{\boldsymbol{B}}\|}{\sqrt{\rho}+\sqrt{\tilde{\rho}}},\quad\ell=1,2,$ (39)
$\alpha_{\ell,r}\left(\boldsymbol{U},\tilde{\boldsymbol{U}}\right):=\max\left\{u_{\ell},\frac{\sqrt{\rho}u_{\ell}+\sqrt{\tilde{\rho}}\tilde{u}_{\ell}}{\sqrt{\rho}+\sqrt{\tilde{\rho}}}\right\}+\mathscr{C}_{\ell}\left(\boldsymbol{U}\right)+\frac{\|\boldsymbol{B}-\tilde{\boldsymbol{B}}\|}{\sqrt{\rho}+\sqrt{\tilde{\rho}}},\quad\ell=1,2,$ (40)
@@
$\mathscr{C}_{\ell}\left(\boldsymbol{U}\right):=\frac{1}{\sqrt{2}}\left[\mathscr{C}_{s}^{2}+\frac{\|\boldsymbol{B}\|^{2}}{\rho}+\sqrt{\left({\mathscr{C}_{s}}^{2}+\frac{\|\boldsymbol{B}\|^{2}}{\rho}\right)^{2}-4\frac{\mathscr{C}_{s}^{2}{B_{\ell}}^{2}}{\rho}}\right]^{\frac{1}{2}}\quad\text{with}\quad\mathscr{C}_{s}=\sqrt{\frac{\left(\gamma-1\right)p}{2\rho}},\quad\ell=1,2.$ (41)
@@ 3.2 Upwind discretization of Godunov–Powell source term
According to the theory in [47, 48], incorporating a properly discretized Godunov–Powell source term into the DG schemes is important to ensure the PP property of the updated cell averages. The source terms can help eliminate the effect of the normal magnetic jump across the cell interface on the PP property. Since the PP property of the updated cell averages is related only to the discrete equations satisfied by the cell averages, we consider adding the Godunov–Powell source term only to the evolution equations of the cell averages. Meanwhile, the evolution of the higher-order “moments” continues to use the OEDG schemes designed in section 2.
Let $\boldsymbol{U}_{\sigma}^{n}\left(x,y\right)$ denote the OEDG solution at the $n$th time level after LDF OE procedure, and let $\bar{\boldsymbol{U}}_{ij}^{n}$ represent its cell average over $I_{ij}$. Since an SSP time discretization can be regarded as a convex combination of the forward Euler method, we only need to discuss the PP property of the OEDG method coupled with the forward Euler time discretization. By incorporating an “upwind” discretized Godunov–Powell source term into the evolution equations for the cell averages, we design
@@
$\boldsymbol{L}_{ij}\left(\boldsymbol{U}_{\sigma}^{n}\right)$ $:=-\frac{1}{\Delta x}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\bigg{[}\left(\widehat{\boldsymbol{F}_{1}}_{,i+\frac{1}{2}}\left(y_{j}^{(\mu)}\right)-\widehat{\boldsymbol{F}_{1}}_{,i-\frac{1}{2}}\left(y_{j}^{(\mu)}\right)\right)$ (45)
$+\mathscr{B}_{1,i+\frac{1}{2}}^{-}\left(y_{j}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{\sigma}^{n}\left(x_{i+\frac{1}{2}}^{-},y_{j}^{(\mu)}\right)\right)+\mathscr{B}_{1,i-\frac{1}{2}}^{+}\left(y_{j}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{\sigma}^{n}\left(x_{i-\frac{1}{2}}^{+},y_{j}^{(\mu)}\right)\right)\bigg{]}$
$-\frac{1}{\Delta y}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\bigg{[}\left(\widehat{\boldsymbol{F}_{2}}_{,j+\frac{1}{2}}\left(x_{i}^{(\mu)}\right)-\widehat{\boldsymbol{F}_{2}}_{,j-\frac{1}{2}}\left(x_{i}^{(\mu)}\right)\right)$
$+\mathscr{B}_{2,j+\frac{1}{2}}^{-}\left(x_{i}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{\sigma}^{n}\left(x_{i}^{(\mu)},y_{j+\frac{1}{2}}^{-}\right)\right)+\mathscr{B}_{2,j-\frac{1}{2}}^{+}\left(x_{i}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{\sigma}^{n}\left(x_{i}^{(\mu)},y_{j-\frac{1}{2}}^{+}\right)\right)\bigg{]}.$
Here the Gauss quadrature nodes $\left\{x_{i}^{(\mu)}\right\}_{\mu=1}^{q}$ and $\left\{y_{j}^{(\mu)}\right\}_{\mu=1}^{q}$ as well as the weights $\left\{\omega_{\mu}^{G}\right\}_{\mu=1}^{q}$ are the same as in (9). The numerical fluxes read
$\widehat{\boldsymbol{F}_{1}}_{,i+\frac{1}{2}}\left(y\right):=\widehat{\boldsymbol{F}_{1}}\left(\boldsymbol{U}_{\sigma}^{n}\left(x_{i+\frac{1}{2}}^{-},y\right),\boldsymbol{U}_{\sigma}^{n}\left(x_{i+\frac{1}{2}}^{+},y\right)\right),$
$\widehat{\boldsymbol{F}_{2}}_{,j+\frac{1}{2}}\left(x\right):=\widehat{\boldsymbol{F}_{2}}\left(\boldsymbol{U}_{\sigma}^{n}\left(x,y_{j+\frac{1}{2}}^{-}\right),\boldsymbol{U}_{\sigma}^{n}\left(x,y_{j+\frac{1}{2}}^{+}\right)\right),$
In (45), we take
$\mathscr{B}_{1,i+\frac{1}{2}}^{\pm}\left(y\right):=\frac{\pm\mathscr{V}_{1,i+\frac{1}{2}}^{\pm}\left(y\right)}{\mathscr{V}_{1,i+\frac{1}{2}}^{+}\left(y\right)-\mathscr{V}_{1,i+\frac{1}{2}}^{-}\left(y\right)}\left[{\left(B_{1}\right)}_{\sigma}^{n}\left(x_{i+\frac{1}{2}}^{+},y\right)-{\left(B_{1}\right)}_{\sigma}^{n}\left(x_{i+\frac{1}{2}}^{-},y\right)\right],$
$\mathscr{B}_{2,j+\frac{1}{2}}^{\pm}\left(x\right):=\frac{\pm\mathscr{V}_{2,j+\frac{1}{2}}^{\pm}\left(x\right)}{\mathscr{V}_{2,j+\frac{1}{2}}^{+}\left(x\right)-\mathscr{V}_{2,j+\frac{1}{2}}^{-}\left(x\right)}\left[{\left(B_{2}\right)}_{\sigma}^{n}\left(x,y_{j+\frac{1}{2}}^{+}\right)-{\left(B_{2}\right)}_{\sigma}^{n}\left(x,y_{j+\frac{1}{2}}^{-}\right)\right]$
with
$\mathscr{V}_{1,i+\frac{1}{2}}^{\pm}\left(y\right):=\mathscr{V}_{1}^{\pm}\left(\boldsymbol{U}_{\sigma}^{n}\left(x_{i+\frac{1}{2}}^{-},y\right),\boldsymbol{U}_{\sigma}^{n}\left(x_{i+\frac{1}{2}}^{+},y\right)\right),$
@@
where $\mathscr{V}_{1}^{\pm}$ and $\mathscr{V}_{2}^{\pm}$ are the wave speeds in the HLL fluxes discussed in section 3.1. Note that
$\mathscr{B}_{1,i+\frac{1}{2}}^{+}\left(y\right)+\mathscr{B}_{1,i+\frac{1}{2}}^{-}\left(y\right)$ $={\left(B_{1}\right)}_{\sigma}^{n}\left(x_{i+\frac{1}{2}}^{+},y\right)-{\left(B_{1}\right)}_{\sigma}^{n}\left(x_{i+\frac{1}{2}}^{-},y\right)=:\left[\!\left[\left(B_{1}\right)_{\sigma}^{n}\right]\!\right]_{i+\frac{1}{2}}(y),$ (46)
$\mathscr{B}_{2,j+\frac{1}{2}}^{+}\left(x\right)+\mathscr{B}_{2,j+\frac{1}{2}}^{-}\left(x\right)$ $={\left(B_{2}\right)}_{\sigma}^{n}\left(x,y_{j+\frac{1}{2}}^{+}\right)-{\left(B_{2}\right)}_{\sigma}^{n}\left(x,y_{j+\frac{1}{2}}^{-}\right)=:\left[\!\left[\left(B_{2}\right)_{\sigma}^{n}\right]\!\right]_{j+\frac{1}{2}}(x)$ (47)
are actually wave-dependent splittings of the jumps in the normal magnetic component across the cell interfaces $x=x_{i+\frac{1}{2}}$ and $y=y_{j+\frac{1}{2}}$, respectively. Such “upwind” splittings are crucial for the PP property of the updated cell averages; see section 3.6.
More specifically, in (45), the terms involving $\mathscr{B}_{1,i+\frac{1}{2}}^{\pm}\left(y\right)$ and $\mathscr{B}_{2,j+\frac{1}{2}}^{\pm}\left(x\right)$ are actually an “unwind” discretization of the cell average of the Godunov–Powell source term
@@
$\frac{1}{\Delta x\Delta y}\int_{y_{j-\frac{1}{2}}}^{y_{j+\frac{1}{2}}}\int_{x_{i-\frac{1}{2}}}^{x_{i+\frac{1}{2}}}\left(-\nabla\cdot\boldsymbol{B}_{\sigma}^{n}\right)\boldsymbol{S}\left(\boldsymbol{U}_{\sigma}^{n}\right)~{}{\rm d}x{\rm d}y.$ (48)
Thanks to the LDF property of $\boldsymbol{B}_{\sigma}^{n}$, the local divergence $\nabla\cdot\boldsymbol{B}_{\sigma}^{n}$ vanishes within each cell $I_{ij}$. Therefore, to approximate (48), we only need to measure the “weak” divergence/derivatives of the normal magnetic component at the cell interface.
Figure 1: The sketch near the interface $x=x_{i+\frac{1}{2}}$: the left/right part of the gray area represents the local region of cell $I_{ij}/I_{i+1,j}$ near the interface with a thickness of $\varepsilon\to 0^{+}$.
Take the interface $x=x_{i+\frac{1}{2}}$ as an example. See Figure 1, where the left/right part of the gray area represents the local region of cell $I_{ij}/I_{i+1,j}$ near the interface $x=x_{i+\frac{1}{2}}$ with a thickness of $\varepsilon\to 0^{+}$. The contribution of the Godunov–Powell source term due to the weak divergence at the interface $x=x_{i+\frac{1}{2}}$ can be approximated by
@@
$\quad\frac{1}{\Delta x\Delta y}\lim_{\varepsilon\to 0^{+}}\int_{x_{i+\frac{1}{2}}-\varepsilon}^{x_{i+\frac{1}{2}}+\varepsilon}\int_{y_{j-\frac{1}{2}}}^{y_{j+\frac{1}{2}}}\left(\frac{\partial\left(B_{1}\right)_{\sigma}^{n}}{\partial x}+\frac{\partial\left(B_{2}\right)_{\sigma}^{n}}{\partial y}\right)\boldsymbol{S}\left(\boldsymbol{U}_{\sigma}^{n}\right)\mathrm{~{}d}x\mathrm{d}y$ (49)
@@
$=\frac{1}{\Delta x\Delta y}\lim_{\varepsilon\to 0^{+}}\int_{x_{i+\frac{1}{2}}-\varepsilon}^{x_{i+\frac{1}{2}}+\varepsilon}\int_{y_{j-\frac{1}{2}}}^{y_{j+\frac{1}{2}}}\frac{\partial\left(B_{1}\right)_{\sigma}^{n}}{\partial x}\boldsymbol{S}\left(\boldsymbol{U}_{\sigma}^{n}\right)\mathrm{~{}d}x\mathrm{d}y$
@@
$\approx\frac{1}{\Delta x}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\lim_{\varepsilon\to 0^{+}}\int_{x_{i+\frac{1}{2}}-\varepsilon}^{x_{i+\frac{1}{2}}+\varepsilon}\frac{\partial\left(B_{1}\right)_{\sigma}^{n}\left(x,y_{j}^{(\mu)}\right)}{\partial x}\boldsymbol{S}\left(\boldsymbol{U}_{\sigma}^{n}\left(x,y_{j}^{(\mu)}\right)\right)\mathrm{~{}d}x$